# A Multi-Stage Auto-Context Deep Learning Framework for Tissue and Nuclei Segmentation and Classification in H&E-Stained Histological Images of Advanced Melanoma

### v1

@@ 2.6 Stage4: Final Tissue segmentation
In this stage, we incorporate the nuclei segmentation result from stage 3 as a fourth input channel and three models similar to those in stage 2 are trained as shown in Figure 4. Tissue ensemble rules used in stage 4 are the same as in stage 2.
Figure 4: The fourth and final stage of the proposed method: Tissue segmentation is refined using nuclei segmentation results. The nuclei segmentation output from the previous stage is added as an additional input channel, and three models similar to those in stage 2 are trained for this task.
@@ 2.7.1 Tissue Post-Processing
Figure 5: Due to the low density of nuclei in necrotic tissue (a), the final segmentation result in stage 4 may suffer from under-segmentation (b). To address this, a post-processing step integrates predictions from stage 1 when they overlap with stage 4 predictions for necrotic tissue (c).
@@ 2.7.2 Nuclei Post-Processing
Figure 6: At image borders, nuclei instance segmentation results are replaced with class map results. (a) Binary segmentation map from HoVer-Next. (b) Binary class map from U-Net ++. (c) Binary map after border correction.
@@ 2.8 Evaluation
To evaluate the performance of the algorithms, the micro Dice score (for multi-class tissue segmentation) and the summed macro F1-score (for nuclei instance detection and classification) were used. The micro Dice score is calculated by treating all segmentation results from all images as a single large image. The summed macro F1-score is determined using a 15-pixel centroid threshold between the predicted segmentation of the nuclei results and the ground truth. For the final ranking of the challenge, the average of the micro Dice score and the summed macro F1-score was used for both Track 1 and Track 2. Further details and implementation of the utilized metrics are available in the PUMA challenge description [^3] https://puma.grand-challenge.org/submission/.
@@ 3 Experiments
To conduct experiments, we utilized both transfer learning (as described in Section2.1) and data augmentation techniques to improve the performance of the models. For data augmentation, we applied both shape-based transformations (flip, rotation, scaling) and intensity-based adjustments (hue, saturation, brightness, contrast). All experiments were conducted using a single workstation with AMD Ryzen 9 7950X CPU, 96 GB of RAM, and two NVIDIA RTX 4090 GPUs. Our implementation for training and testing is available on our GitHub repository: https://github.com/NimaTorbati/PumaSubmit.
@@ 4 Results & Discussion
As the results show, all three top-ranked teams outperformed the provided baseline method by a large margin. Details of the methods used by the other top-performing teams (i.e., TIAKong, rictoo, and agaldran) are not available at the time of writing this manuscript but are expected to be released by the organizers. With our proposed approach (LSM), we achieved second place in Track 2 and first place in Track 1 of the competition.
When considering only the performance for nuclei instance segmentation and classification, we achieved the best performance in Track 1 and the second-best performance in Track 2. However, for multi-class tissue segmentation, our method ranked third in Track 1 and second in Track 2.
Table 1: The final test results and ranking of Track 1 of the PUMA challenge.
@@
Table 2: The final test results and ranking of Track 2 of the PUMA challenge.
@@
As pointed out in Section 2.7, we only applied the proposed tissue segmentation post processing in Track 2 of the challenge. As the average results indicate, this step significantly improved overall performance. We also report the tissue-specific scores on the final test sets of Track 1 and Track 2 in Table 3. According to these results, the post processing step notably enhanced the segmentation performance for necrosis tissue (from 46.69% to 74.49%). However, it should be noted that necrosis is the least represented tissue across the training, validation, and test sets. Therefore, even small over- or under-segmentation can significantly affect the micro Dice score for necrosis, and thus influence the overall average tissue segmentation performance.
In Table 3, we also compare our tissue-specific performance with that of the top-performing team (TIAKong). As the results show, our method outperforms TIAKong for epidermis and blood vessel tissue segmentation, while TIAKong delivers better results for the remaining tissue types on the final test set.
@@
It should be noted that the summed F1-score for each nuclei type is not provided by the challenge organizers; only the average results are reported as shown in Table 1 and Table 2. When this information becomes available, we will include it in the updated version of the manuscript. However, in Figure 7, we show an example result of our approach and compare it with the baseline HoVer-Next and the ground truth segmentation mask. As the qualitative results clearly show, our predicted segmentation mask is much closer to the ground truth compared to the baseline HoVer-Next.
@@
Figure 7: Nuclei segmentation and classification result example
Further detailed results, including our internal and external validation as well as findings from the ablation study demonstrating the impact of each stage and the tissue and nuclei post processing steps, will also be added to the updated version of this manuscript in the near future.
@@ 5 Conclusion
In this work, we introduced a novel multi-stage auto-context deep learning framework to perform tissue and nuclei segmentation and classification in H&E-stained histological images of advanced melanoma. By exploiting contextual information across stages and integrating multiple state-of-the-art segmentation models, our method achieved excellent performance in the PUMA Challenge, ranking second in Track 1 and first in Track 2 of the competition. These results confirm the effectiveness of combining tissue and nuclei information in a unified framework. In future work, we plan to evaluate the performance of the proposed method on an external dataset and report the impact of each stage on the final performance through a comprehensive ablation study.

### v2

@@ 2.6 Stage 4: Final Tissue Segmentation
In this stage, we incorporate the nuclei segmentation result from stage 3 as a fourth input channel to train two SegFormer models similar to those in stage 2 as shown in Figure 4. We used the same U-Net trained model in stage 2 without incorporating the fourth nuclear channel, in order to simplify the framework while not degrading the overall performance, particularly for blood vessel detection. Tissue ensemble rules used in stage 4 are the same as in stage 2.
Figure 4: The fourth and final stage of the proposed method. Tissue segmentation is refined using nuclei segmentation results. The nuclei segmentation output from the previous stage is added as an additional input channel, and two SegFormer models similar to those in stage 2 are trained. For the U-Net model, we used the same trained model in stage 2 without incorporating the fourth nuclear channel. Color code for tissue segmentation: red for tumor, blue for stroma, and green for blood vessel; color code for nuclei segmentation: various cell types are indicated by different colors.
@@ 2.7.1 Tissue Post-Processing
Figure 5: Post-processing of necrotic tissue. Due to the lower density of nuclei in necrotic part of the tissue (a), the final segmentation result in stage 4 may suffer from under-segmentation (b). To address this, a post-processing step integrates predictions from stage 1 when they overlap with stage 4 predictions for necrotic tissue (c). Color code in (b) and (c): yellow for necrotic part of the tissue, red for tumor.
@@ 2.7.2 Nuclei Post-Processing
Figure 6: Post-processing at image borders. At image borders, nuclei instance segmentation results are replaced with class map results. (a) Binary segmentation map from HoVer-Next. (b) Binary class map from U-Net ++. (c) Binary map after border correction. Color code: white for nuclei, black as background. Exemplified differences are shown in red boxes.
2.8 External Validation on the MoNuSAC Dataset
The MoNuSAC dataset contains only nuclei annotations and does not include tissue-level annotations. Therefore, we adapted the proposed framework for this dataset as follows. Since skin tissue images are not present in the MoNuSAC data, Stage 1 was omitted, and all images were treated as one class and models trained with metastatic images were used. In Stage 2, we applied the pre-trained tissue segmentation model from the PUMA dataset to generate tissue predictions. These predictions were then used in Stage 3, where the nuclei segmentation model was trained using the MoNuSAC training images. For the instance segmentation branch in Stage 3, we separately used nuclei instance masks provided by the MoNuSAC challenge winner team [26] and the post-challenge winner team [30], both of which are publicly available on the MoNuSAC challenge website. The performance of the proposed method was evaluated on the test images of the MoNuSAC dataset.
@@ 2.9 Evaluation
To evaluate the performance of the algorithms, the micro Dice score for multi-class tissue segmentation and the summed macro F1-score for nuclei instance detection and classification were used for the PUMA dataset as the official evaluation metrics of the challenge. The micro Dice score is calculated by treating all segmentation results from all images as a single large image. The summed macro F1-score is determined using a 15-pixel centroid threshold between the predicted segmentation of the nuclei results and the ground truth. For the final ranking of the PUMA challenge, the average of the micro Dice score and the summed macro F1-score was used for both Track 1 and Track 2. Further details and implementation of the utilized metrics are available in the PUMA challenge description [^8] https://puma.grand-challenge.org/submission/.
For the MoNuSAC dataset, we adopted the class-specific Panoptic Quality (PQ) score, consistent with the official MoNuSAC challenge evaluation protocol. Additionally, inspired by the micro Dice score, we calculated the micro PQ by treating all segmentation results from all images as a single large image. Further details and implementation of the PQ are available in the MoNuSAC challenge description [^9] https://monusac-2020.grand-challenge.org/Evaluation_Metric/.
@@ 3 Experiments
To conduct experiments, we utilized both pre-training with the PanopTILs dataset (as described in Section 2.1) and data augmentation techniques to improve the performance of the models. For data augmentation, we applied both shape-based transformations (flip, rotation, scaling) and intensity-based adjustments (hue, saturation, brightness, contrast). All experiments were conducted using a single workstation with AMD Ryzen 9 7950X CPU, 96 GB of RAM, and two NVIDIA RTX 4090 GPUs. Our implementation for training and testing is available on our GitHub repository: https://github.com/NimaTorbati/PumaSubmit.
@@ 4 Results & Discussion
This section presents the evaluation results of the proposed method on the PUMA and MoNuSAC datasets. Additionally, it includes an ablation study analyzing the contribution of each stage within the proposed framework.
4.1 PUMA Dataset Results
@@
As the results show, all three top-ranked teams outperformed the provided baseline method by a large margin. Details of the methods used by the other top-performing teams among those TIAKong, rictoo, and agaldran are not available at the time of writing this manuscript but are expected to be released by the organizers on the challenge website. With our proposed approach (LSM), we achieved second place in Track 2 and first place in Track 1 of the competition.
When considering only the performance on nuclei instance segmentation and classification, we achieved the second-best result in Track 1 (behind the rictoo team with 75.77%) and the best performance in Track 2. However, for multi-class tissue segmentation, our method ranked third in Track 1 (following TIAKong with 78.23% and Biototem with 72.68%) and second in Track 2 (after TIAKong with 78.23%).
Table 1: The final test results and ranking of Track 1 of the PUMA challenge (%).
@@
Table 2: The final test results and ranking of Track 2 of the PUMA challenge (%).
@@
As pointed out in Section 2.7, we only applied the proposed tissue segmentation post processing in Track 2 of the challenge. As the average results indicate, this step significantly improved overall performance.
Another observation from the results in Table 1 and Table 2 is the significant performance drop in nuclei classification and segmentation as the number of nuclei classes increases. For example, the summed macro F1 score for Track 1 (with three nuclei classes) using our proposed approach is 74.43%, whereas for Track 2 (with ten nuclei classes), it drops to 48.97%. This performance decrease could be attributed to the increased complexity of Track 2 compared to Track 1. Additionally, it may be related to the challenges of accurately annotating H&E-stained images when a larger number of nuclei classes are involved. This limitation also has biological basis and is biologically relevant, as other staining techniques—such as immunohistochemistry (IHC)—are commonly applied to visualize individual cell types using specific cellular markers in histological images. Consequently, detection and classification of certain nuclei types based solely on H&E staining may not be feasible. To address this challenge and achieve accurate nuclei segmentation and classification for specific cell types, using specially designed and annotated datasets based on marker-specific staining techniques could be a promising solution.
We also report the tissue-specific scores on the final test sets of Track 1 and Track 2 in Table 3. According to these results, the post processing step notably enhanced the segmentation performance for necrosis tissue (from 46.69% to 74.49%). However, it should be noted that necrotic part of tissue is the least represented tissue class across the training, validation, and test sets. Therefore, even small over- or under-segmentation can significantly affect the micro Dice score for necrosis, and thus influence the overall average tissue segmentation performance. In Table 3, we also compare our tissue-specific performance with that of the top-performing team (TIAKong). As the results show, our method outperforms TIAKong for epidermis and blood vessel segmentation, while TIAKong delivers better results for the remaining tissue types on the final test set.
In Table 4 and Table 5, the nuclei-specific performance of our proposed approach, as well as that of the top-performing team (TIAKong) and the baseline model, is presented. As shown in Table 4 for Track 1, which includes three nuclei classes, our approach delivers superior performance for the Other class, while TIAKong achieves slightly better results for the Tumor and Lymphocyte classes. Compared to the baseline model, our approach outperforms it across all three nuclei classes. As depicted in Table 5 for Track 2, our method outperforms TIAKong for apoptosis, epithelium, histiocyte, neutrophil, and stroma, while TIAKong achieves better performance for the remaining nuclei types on the final test set. Once again, compared to the baseline model, our approach delivers superior performance across all ten classes by a large margin.
@@
Table 4: Macro F1 scores of Track 1 for tumor, lymphocyte, and other nuclei classes (%).
Team Tumor Lymphocyte Other Macro F1
LSM 82.87 79.24 61.17 74.43
TIAKong 83.86 81.04 58.26 74.39
Baseline 77.68 69.40 51.25 69.40
Table 5: Macro F1 scores of Track 2 for each nuclei class (%).
Class LSM TIAKong Baseline
Apoptosis 38.00 19.46 5.86
Endothelium 43.93 44.03 23.92
Epithelium 75.35 60.83 3.07
Histiocyte 45.66 43.22 36.11
Lymphocyte 75.34 80.90 69.81
Melaphonage 38.00 44.48 29.52
Neutrophil 29.39 26.97 16.96
Plasma 21.14 25.00 10.15
Stroma 40.56 38.13 29.07
Tumor 82.29 83.83 73.20
Macro F1 48.96 46.69 29.77
For visualization of nuclei segmentation, Figure 7 presents a sample result from our method compared with the baseline HoVer-NeXt and the ground truth segmentation mask. As the qualitative results demonstrate, our predicted segmentation is notably closer to the ground truth than the baseline.
For tissue segmentation, Figure 8 presents an example result from our method alongside the baseline nnU-Net for comparison, where our model delivers qualitatively superior results.
@@
Figure 7: Example of nuclei segmentation and classification results from the PUMA dataset. GT: Ground Truth.
Figure 8: Example of tissue segmentation and classification results from the PUMA dataset. GT: Ground Truth.
4.2 MoNuSAC Dataset Results
As explained in Section 2.8, we used the instance nuclei segmentation results from the MoNuSAC challenge winning team (TIA-Lab) and the post-challenge winning team (AmirrezaMahbod or AM) to investigate the impact of integrating the proposed multi-stage and auto-context approach on their performance. The results of these comparisons are presented in Table 6. As the nuclei-specific and average nuclei scores (mean Micro PQ and mean PQ) indicate, our proposed approach yields consistent improvements across all metrics for both baseline methods, with the exception of the macrophage class in the AM approach. These improvements suggest that leveraging global tissue-level information enhances the model’s ability to distinguish between different nuclei types, particularly in complex histological environments. It is worth mentioning that, for generating tissue segmentation masks in the MoNuSAC dataset, we used the PUMA tissue segmentation pre-trained model without any additional annotations or training from the MoNuSAC dataset.
Table 6: Comparison of class-specific and average micro Panoptic Quality (PQ) scores, as well as mean PQ, on the MoNuSAC test set.
Method Epithelial Lymphocyte Neutrophil Macrophage Mean Micro PQ Mean PQ
TIA-Lab 70.52 70.04 68.02 37.52 61.52 61.04
TIA-Lab with LSM 71.15 70.54 69.86 39.26 62.71 62.29
AM 67.23 68.63 62.74 52.58 62.82 59.57
AM with LSM 68.71 69.53 68.01 50.74 64.22 62.18
An example result from the MoNuSAC dataset is illustrated in Figure 9. The inclusion of tissue-level information improves the performance of nuclei classification. As shown in the figure, falsely recognized lymphocytes in TIA-Lab and AM methods are correctly classified as epithelial cells when LSM method is additionally integrated.
Figure 9: Example nuclei segmentation and classification results from the MoNuSAC dataset. The figure compares ground truth annotations with predictions from the MoNuSAC challenge winner (TIA-Lab), the post-challenge winner (AM), and the proposed method (LSM). The proposed method benefits from tissue context integration, resulting in improved nuclei type classification.
4.3 Ablation Study
To better understand the contribution of individual stages to the overall performance of the proposed framework, we conducted an ablation study by systematically removing key components of the model. Specifically, we selected six components and designed seven experimental configurations by excluding these elements. All ablation experiments were performed using the internal validation set to assess the performance impact of each component.
Experiments 1 and 2: Effect of the Classifier
To evaluate the impact of the classifier, we removed Stage 2 entirely and directly used the tissue segmentation output from the SegFormer model in Stage 1. As shown in the Exp1 and Exp2 rows of Table 7, this modification resulted in decreased segmentation performance for both tissue and nuclei. These results highlight the importance of incorporating a dedicated classifier, especially for tasks involving small datasets.
Experiment 3: Effect of the Classification Rules
As discussed in Stage 1, two classification rules are used: (1) if Epidermis is present in the segmentation, the image is classified as primary, and (2) if the number of primary pixels exceeds metastatic pixels, the image is classified as primary. In Experiment 2, only the second rule was applied, whereas in Experiment 3, both rules were used. As shown in Table 7, incorporating the first rule led to a slight improvement in tissue segmentation performance.
Experiment 4: Effect of the U-Net Branch
The U-Net model was integrated into both Stage 2 and Stage 4 to enhance blood vessel tissue segmentation. Our findings indicate that U-Net outperforms SegFormer for detecting blood vessels. Incorporating U-Net improved segmentation performance for both tissue and nuclei (Table 7). Given that blood vessels are small structures, these results suggest that SegFormer performs better on classes with larger spatial dependencies compared to U-Net.
Experiment 5: Effect of Stage 4 (Tissue Refinement Using Nuclei Information)
In Stage 4, nuclei segmentation results are used to refine tissue predictions. In Experiment 5, we included this refinement step to assess its impact. The results demonstrate a clear improvement in overall tissue segmentation performance (Table 7). This supports the effectiveness of an auto-context strategy, where the model leverages its own outputs to enhance performance.
Experiments 6 and 7: Effect of Ensemble Rules and Post-processing
Finally, we evaluated the impact of the ensemble rules and post-processing techniques. As shown in Table 7, these modifications led to a slight improvement in tissue segmentation accuracy, further refining the performance of the proposed method. It is worth noting that, compared to the results on the final test set of the PUMA challenge, the effect of post-processing on the internal validation dataset has a smaller impact.
Table 7: Ablation study results showing the effect of each module on segmentation performance. The last row (Mean) is the average of Macro F1, Dice, and Micro Dice. Best values in each metric are highlighted in bold.
Component Exp1 Exp2 Exp3 Exp4 Exp5 Exp6 Exp7
Classifier
Classification Rules
Unet Branch
Stage 4
Tissue Ensemble Rules
Post Processing
Macro F1 31.47 32.26 32.26 33.88 33.88 33.88 33.88
Dice 74.06 81.08 81.09 80.04 82.05 81.93 82.11
Micro Dice 63.44 72.99 73.00 74.91 77.65 77.76 77.66
Mean 56.32 62.11 62.11 62.94 64.52 64.52 64.55
@@ 5 Conclusion
In this work, we introduced a novel multi-stage auto-context deep learning framework for tissue and nuclei segmentation and classification in H&E-stained histological images of advanced melanoma. By leveraging contextual information across stages and integrating multiple state-of-the-art segmentation models, our method achieved excellent performance in the PUMA Challenge, ranking second in Track 1 and first in Track 2. Furthermore, the proposed approach outperformed the winners of the MoNuSAC challenge in terms of nuclei classification. These results highlight the effectiveness of fusing tissue and nuclei information within a unified framework.